# A Population Analysis of 20 Exoplanets Observed from the Optical to the Near-infrared Wavelengths with HST: Evidence for Widespread Stellar Contamination

### v2

@@ 1 Introduction
We recently surpassed the mark of 5600 confirmed exoplanets, with more planets discovered by the day thanks to different and complementary detection techniques. Studies involving large populations of planets and their stellar hosts have been pivotal to get a glimpse into the great diversity of planetary systems beyond our own. Notable examples of said studies focused on exoplanet demographics include e.g. Fulton et al. (2017); Howard et al. (2012); Mayor et al. (2011); Cassan et al. (2012); Nielsen et al. (2019); Gaia Collaboration et al. (2022); Luque & Pallé (2022). Exoplanet detection campaigns and dedicated missions, such as NASA Kepler, TESS and ESA Cheops, have also significantly contributed to characterise their host stars by measuring key stellar parameters and monitoring their magnetic activity, see e.g. Bastien et al. (2013); Yang et al. (2017); Basri et al. (2013); Balona (2015, 2022); Magrini et al. (2022); Simpson et al. (2023).
Large-scale spectroscopic studies of exoplanetary atmospheres are today in their infancy. Space telescopes, such as NASA Spitzer and NASA/ESA Hubble, have delivered throughout their years of operation large enough data-sets to enable pioneering population studies for tens of exoplanet atmospheres, mainly hydrogen-rich (see e.g., Cowan & Agol (2011); Iyer et al. (2016); Sing et al. (2016); Fu et al. (2017); Tsiaras et al. (2018); Fisher & Heng (2018); Changeat et al. (2022); Edwards et al. (2022)). However the data quality provided by these space telescopes did not always match the ambition of the scientific community. To enable detailed comparative exoplanetology, population studies should ideally satisfy the following conditions:
Exoplanet spectra should be simultaneously recorded across the broadest possible wavelength range to mitigate the hindrance of observations taken at different epochs (Danielski et al., 2014). This issue is particularly acute if the stellar host is active and/or the atmosphere shows signs of temporal variability due to changes in the cloud coverage, climate or weather patterns (Armstrong et al., 2016; Cohen et al., 2023; Changeat et al., 2024). In particular, optical-to-infrared atmospheric spectra may help to address degeneracies among cloud parameters (Mai & Line, 2019) and to correct for stellar contamination (Ballerini et al., 2012). These mitigations allow to constrain better chemical abundances (Rustamkulov et al., 2023) and the thermal structure of the atmosphere.
@@
While significant strides have been made in characterising exoplanetary atmospheres, less emphasis has been placed so far on using the spectral information in the optical/near-UV regime to detect and correct for potential stellar contamination (e.g. Pinhas et al., 2018; Edwards et al., 2021; Libby-Roberts et al., 2022). The difficulty of this task arises from the diverse ways in which an active star can influence the spectra of its companion planet(s). Stellar magnetic activity results in surface heterogeneities in the form of colder spots and hotter faculae that can cause the transit chord to differ from the observed disk-integrated stellar spectrum. On the stellar photosphere, these heterogeneities can either be occulted or unocculted during the planetary transit w.r.t. our line of sight. While occulted stellar signatures are comparatively easier to identify in transit light curves, unocculted features are trickier to recognise. The first imprint a bump (in the case of a spot) or a depression (for a facula) in the transit light curve, but can be masked or removed by inspection (Narita et al., 2013). On the other hand, unocculted active regions affect the transit depth of the entire light curve, introducing a strong chromatic signal that deepens (decreases) the observed transit depth, overestimating (underestimating) the planetary radius compared to a scenario without spots (faculae) (Rackham et al., 2018). Although stellar contamination effects are strongest in the optical regime, where they introduce characteristic, exponential slopes, they are also non-negligible in the infrared where they are capable of introducing offsets, potentially mimicking or masking absorption features of interest in the planetary atmosphere, e.g. water (Barclay et al., 2021) and/or resulting in erroneous retrieved chemical abundances (Iyer & Line, 2020; Thompson et al., 2024).
@@ 2 Data Analysis
We searched the MAST archive for publicly available exoplanet transit data observed, with at least one grating/grism, by both the STIS and WFC3 instruments on HST. Although we were able to identify more than 30 planets that met this criteria, the data sets considered for this study refer to 20 planets only. To produce a consistent analysis and have a homogeneous data sample, we discarded HAT-P-70 b and KELT-20 b from the final selection as their STIS observations were taken with STIS/G430 at medium resolution. Similarly, we did not include WASP-43 b as its G430L data set did not display a transit and its G750 observation was taken with the medium resolution grating (G750M). WASP-12 b, despite being observed by both grisms on STIS and WFC3, was excluded from our sample due to the contamination by a nearby companion within the system. To obtain a reliable spectrum of WASP-12 b, meticulous decontamination is essential, involving modelling of the companion star. Planetary spectra that displayed strong offsets among their STIS and WFC3 observations were also rejected. These include: HAT-P-17 b, HAT-P-32 b, HAT-P-41 b, HD 189733 b, KELT-7 b, WASP-19 b, WASP-62 b, WASP-76 b, WASP-79 b. Extra processing steps are necessary to correct for the observed offsets and to explain their origin, therefore these planets will be the subject of a separate study. The final sample selection is displayed in Figure 1, while the observations’ proposal ID and the names of their respective principal investigators are reported in Appendix A. There we also include the instrument-filter combination each planet was observed with and the date each observation was performed. We are aware that some of these exoplanets have also been observed with other instruments, including JWST, but for homogeneity reasons, here we restrict our analysis to data acquired with the two low resolution gratings on STIS and the infrared channel of WFC3. For the avoidance of doubt, each transit spectrum presented here was taken with all or a combination of the G430L and G750L grating on STIS, and the G102 and G141 grism on WFC3, specifically in spatial scanning mode.
@@
While most of the STIS data sets we utilise have undergone prior analysis by different groups employing diverse data reduction pipelines, our study marks the first instance in the literature of analysing the G430L and G750L data sets of HAT-P-18 b and the G430L data set of HAT-P-26 b. In our wide-sample investigation, we introduce a novel data reduction framework for STIS data, previously presented in Saba et al. (2022), that enables us to uniformly process numerous data sets, with the goal of mitigating biases. WFC3-IR spatial scanned observations were also treated homogeneously across the target sample either by including in our study data sets previously analysed with the Iraclis pipeline or, where new observations were made available, by reducing them with Iraclis (Tsiaras et al., 2016a). Such approach guarantees a standardised methodology for the selected sample and enables a comparative study.
@@ 2.2 WFC3 Data Processing
To ensure homogeneity in the data analysis process, we employed one single pipeline to analyse all transit data sets. Iraclis (Tsiaras et al., 2016a), a specialised WFC3 spatial scanning mode data reduction pipeline and recently adapted to analyse STIS data as well, was used in this study. Most of the WFC3 data sets presented here were previously analysed with Iraclis and collected in two major population studies by Tsiaras et al. (2018) and Edwards et al. (2022). However, these studies focused predominately on WFC3/G141 data, while here we aim to include WFC3/G102 as well. For this reason, we downloaded all publicly available WFC3/G102 data sets for the planets in our sample and processed them with Iraclis, after ensuring that the planetary system parameters (Appendix B) we employed were consistent with those used in previous WFC3/G141 studies and across the remaining data sets associated with the same planet. The planets observed with G102 in transmission include HAT-P-11 b, HAT-P-26 b, WASP-17 b and WASP-39 b. These are also the planets that possess a complete spectrum from 0.4 to 1.6 µm, i.e. were observed with STIS/G430L, STIS/G750L, WFC3/G102 and WFC3/G141, albeit at different epochs.
@@ 2.3 Light curve fitting
Similarly to WFC3, we attempted to perform the detrending of the STIS spectral light curves with the divide-white method (Kreidberg et al., 2015) but, likely due to chromatic effects arising from the host star and the well-known wavelength-dependent thermal breathing of the STIS detector (Brown et al., 2001; Sing et al., 2011), this approach was unsuccessful. A variety of spectroscopic STIS studies have highlighted how the systematics can be both time-dependent (e.g. Nikolov et al., 2014; Lothringer et al., 2018) and wavelength-dependent during a single observation (Huitson et al., 2013), therefore needing a specialised treatment case by case. However, we opted for a uniform systematics model throughout all data sets to maintain consistency: to correct for the combined instrumental and astrophysical trends characteristic of the STIS spectral light curves, we adopted the usual systematics model in Equation 2 and applied it to each spectral light curve separately. Effectively, we modelled the spectral light curves in the same manner as the white light curve. This approach has been used extensively in past STIS studies and has been shown to perform better than employing the divide-white technique (e.g. Chachan et al., 2019). It has consistently proven effective in detrending all data within our sample, thereby ensuring reliable and comparable results throughout.
In the literature there are many examples of how Gaussian Processes (GPs) are employed to fit exoplanet transit light curves (Gordon et al., 2020), especially at the optical wavelengths (Gibson et al., 2017). Generally these algorithms are able to smooth out external features that do not belong to the phenomenon intended to be modelled, in this case by removing any potential spot-crossing signatures from the transit signal. Hence, we attempted to utilise them to detrend STIS light curves as reported in Appendix D. However, we found GPs to be ineffective in providing additional constraints on the light curve modelling. Therefore, all spectral data presented here and employed during the retrievals were produced by applying the standard linear plus exponential detrending as described in Equation 2.
@@ 2.4 Parameters estimation: and
During white light curve fitting we fixed the and parameter values to those obtained from literature. Because of the gaps in the light curves, HST observations are not suited to constrain these parameters. A significant coverage in ingress and/or egress is necessary to constrain the semi-major axis and orbital inclination. To test whether combining many HST light curves could be a suitable approach to derive these parameters in the absence of complete light curves, we compared the values obtained by combining three STIS white light curves with those obtained by fitting a complete JWST/NIRISS light curve. We selected WASP-39 b as our candidate planet, as at the time of writing this is the only planet in our sample for which there exist public JWST observations. We used again Equation 2 to detrend the light curves, but leaving and as free parameters. We compared the results obtained using these two approaches with the values listed in ExoClock, and we conclude that the fitting of JWST light curves leads to the most constrained results, which are compatible well within 1 of the ExoClock parameters. The and values obtained from the combined STIS light curve fitting ( = , = ) have a ten times larger uncertainty compared to the JWST fitted values ( = 11.4190.026, = 87.7720.032, = 0.44390.0062) but comparable to the ExoClock uncertainties ( = , = ). The semi-major axis and inclination parameters derived from the fitting of the three STIS light curves are 2 and 3 away respectively from both JWST and ExoClock, while JWST and ExoClock parameters agree within 1. Therefore, rather than trying to constrain these orbital parameters by fitting multiple incomplete light curves together, we suggest employing either tabulated parameters or fitted parameters from complete white light curves and fixing them during the spectral light curve fitting. For consistency we fitted all white and chromatic light curves using the ExoClock parameters as fixed values across the whole planetary sample, including on WASP-39 b.
@@ 3 Spectral Modelling
Due to the often divergent STIS spectra resulting from their acquisition at varying epochs (Fig. 2), we decided to model the different combinations of data sets as separate spectral scenarios for each planet. As reported in Table 1, each of the potential combinations of data sets is identified by a Case number. The spectral modelling is performed with the fully-Bayesian retrieval framework TauREx 3.1 (Al-Refaie et al., 2021, 2022). Through the use of nested sampling techniques (Skilling, 2006), the retrieval code maps the correlations among the atmospheric parameters in the defined parameter space. TauREx operates on an entirely open-source framework. Utilising the plug-in system, users have the flexibility to integrate their own models, allowing for the representation of various atmospheric conditions such as stellar contamination, cloud parameterisation, and temperature-pressure profiles of their choosing. To explore the parameter space, we employed the widely utilised nested sampling algorithm MultiNest (Feroz et al., 2009; Buchner et al., 2014) with 250 live points and an evidence tolerance set to 0.5.
3.1 Retrievals Setup
[…]
where is the host star’s effective temperature, is the stellar radius, is the inverse of the planet’s semi-major axis to stellar radius ratio. When calculating , we set the albedo () and the emissivity () to be 0.2 and 0.8 respectively. The error bars associated with the equilibrium temperatures presented in Table B.1 were computed under the assumption of different albedo and emissivity values. For the lower bound we adopted and , indicating a planet with high reflectivity and radiative properties. For the upper bound, assuming a condition with low reflectivity but a high heat redistribution potentially resulting from the planet’s formation process, we employed and . In all of our retrievals, we modelled the stellar photosphere using the PHOENIX BT-Settl synthetic stellar atmosphere spectra (Allard et al., 2012). We fixed the fundamental stellar parameters, the stellar radii () and masses (), photospheric effective temperatures () and metallicities to literature values, as reported in Table C.1. To account for potential stellar contamination within the retrieval, four additional parameters relating to stellar activity are fit for alongside the planetary parameters of interest. These activity parameters are the spot and facula temperatures, and respectively (assuming that all spots/faculae are characterised by a single temperature), and their corresponding filling factors i.e. coverage as a fraction of the total projected stellar surface, which are denoted as and . The prior bounds on the filling factors and were set between 0.0 and 0.9, where a filling factor of 0.9 would equate to 90% of the photosphere being covered by either of these active regions, with the implicit condition that the sum of these two parameters cannot exceed 1. We set prior bounds for the temperatures of these active regions that are dependent on the effective (photospheric) temperature () of the host star and, as such, their absolute values vary on a case-by-case basis. For the spot temperature, , we set the lower bound to be equal to - 1500K and the upper bound to be equal to . For the facula temperature, , we set the lower bound to be equal to and the upper bound as + 500K. The reasoning behind imposing priors of different magnitudes on the faculae and spot temperatures is that we expect faculae to display a smaller positive temperature contrast () with respect to the photosphere as compared to the larger negative contrast of spots (e.g. Berdyugina, 2005; Gondoin, 2008; Panja et al., 2020; Norris et al., 2023).
4 Stellar Activity metrics
4.1 Bayes Factor
To assess the strength of the model preference from a Bayesian perspective, we utilise the Bayes Factor, defined as the ratio of two models’ evidences given an observed data set. In our study we compare the evidence of the baseline model to that of the stellar baseline model for each planetary spectra, considering one spectral case at a time. The fitted parameters for the base model include water vapour, , and , while the stellar model incorporates the four additional activity parameters described in Section 3.1 above, thus increasing the dimensionality by four. All models used in the Bayes Factor calculations are cloud-free and assume a uniform atmospheric temperature and water content with altitude. By employing the simplest atmospheric models we aim to distinguish at first order the signatures that stellar activity imprints on the planetary spectra. The Bayes Factor serves as a fully Bayesian approach to model selection, relying solely on the evidence of the respective models. In log space it is expressed as follows:
where represents the Bayesian evidence of the stellar model and corresponds to that of the base model. A negative value suggests a preference for the base model, while a positive value indicates a preference for the model incorporating stellar activity, despite the additional dimensionality it requires. Following the Jeffreys’ scale (Jeffreys, 1998; Trotta, 2008) we establish three thresholds to gauge the extent to which the stellar model is favoured over the baseline model (or vice versa). A Bayes factor 1 2.5 represents a weak evidence against the simpler base model; 2.5 5 is considered to be moderate evidence in favour of including stellar contamination and 5 is considered strong evidence in favour of using a model that incorporates stellar activity. Similarly, the Jeffreys scale can be used to explain the preference of the base model over the stellar model when the Bayes Factor is negative. For instance, a -5 signifies a strong preference for the base model over the active star model. A Bayes factor falling between -5 and -2.5 suggests a moderate preference for the baseline model, while a ranging from -2.5 to -1 indicates weak evidence for it. The strength of the evidence becomes inconclusive for values 1 -1, effectively implying that both models are equally capable of explaining the observed data. The Bayes Factor is a robust indicator for goodness-of-fit and an essential criterion to discriminate against models. Arguably, the greatest strength of the Bayes factor is its tendency to favour simpler models over complex ones in order to prevent over-fitting. This is essential to make sure that our retrieval model is not biased towards finding stellar activity if there is not sufficient evidence for it. However, due to how Bayesian frameworks will naturally interact with low resolution spectra and their associated uncertainties, to obtain a bigger picture perspective of the activity environment we also explore additional, newly defined metrics to explore the contexts in which stellar activity plays a relevant role and discern its extent on a qualitative scale, e.g. low to high activity regimes. A complete overview of these metrics will be given in Thompson et al. (in prep.), where a more extensive description will be presented alongside further applications.
4.2 Stellar Activity Distance metric (SAD)
The Bayesian approach to model selection tends to favour data sets with smaller uncertainties, prioritising precise data during the fitting routine. Consequently, STIS observations, which typically exhibit larger error bars compared to WFC3 due to their lower S/N, will carry less weight during the retrieval process, leading the model to assign less significance to the STIS data points. However, since stellar activity tends to imprint stronger signals within the wavelength range covered by STIS, it is crucial to evaluate how well the retrieval aligns with observations in the visible spectrum and to extract as much of the information content of this region as we can. Solely relying on STIS data sets for retrievals is not advisable; meaningful results require anchoring the optical observations to WFC3 data where the bulk of the planetary information content lies. Without the continuum provided by the near-infrared data, the task of constraining basic planetary parameters becomes exceedingly challenging, resulting in larger uncertainties, increased model and parameter degeneracies and a tendency for the retrieval to be pushed towards exploring unphysical territories. Hence, we introduce a metric that, focusing exclusively on the optical regime, aims at evaluating which retrieval model (with or without activity) better aligns with STIS observations without the need for running further retrievals or altering the input observation in any way.
with an uncertainty of
4.3 Stellar Activity Temporal metric (SAT)
To assess the repeatability and consistency of observations on each planet, we examine how the data sets acquired with the same STIS grating compare to each other. We note that stellar activity is not the only phenomenon which could be responsible for spectra taken at different epochs diverging from one another, other temporally-variable processes e.g. heterogeneous, patchy clouds (MacDonald & Madhusudhan, 2017) or the influence of macro-scale planetary weather/storms (Skinner & Cho, 2022) could also contribute. However, stellar activity is likely the most plausible cause of such strong deviations observed for some of the planets in our sample over these relatively short timescales due to its modulation preliminarly on the timescales of stellar rotation. This is especially true of observations where we see a reversal in the optical slope between subsequent STIS visits as seen for e.g. WASP-6 b, WASP-17 b, WASP-80 b and HD 209458 b which may represent the host star transitioning between spot-dominated and faculae-dominated regimes. This metric cannot be used for WASP-127 b and HAT-P-18 b since we possess only one observation per grating for each of them. In this context, we disregard retrieval models and concentrate purely on observational data. Our aim is to compare two or more observations of the same planet normalised to a common factor. To maintain uniformity across the entire sample, we use the WFC3 data set of each planet as the normalisation factor for all its STIS observations considered for this metric. For each planet, the normalisation factor is determined as the weighted mean of its WFC3 data set, with weights assigned inversely proportional to the transit depth error bars, thereby giving less weight to larger error bars. Finally, we compute the average of the set of absolute normalised differences per pair of observations ( and ), which we define as the SAT:
For individual planets, determining the average deviation in transit depth between subsequent STIS visits in ppm would be informative on its own, however, to perform comparative planetology we need a method of normalising these differences across our entire population. As each of these planets have atmospheres characterised by different scale heights, for comparability we report the absolute differences as a percentage of the weighted average transit depth of the WFC3 observations, which are available for every planet in our sample. Our assumption here is that at first order, the weighted average transit depth taken from the WFC3 observation is a reasonable proxy for scale height effects, although it will of course be slightly influenced by any absorption features present in that region.
5 Results
5.1 Retrieved Temperature and Water Abundance
[…]

### v3

@@ 1 Introduction
We recently surpassed the mark of 5700 confirmed exoplanets, with more planets discovered by the day thanks to different and complementary detection techniques. Studies involving large populations of planets and their stellar hosts have been pivotal to get a glimpse into the great diversity of planetary systems beyond our own. Notable examples of said studies focused on exoplanet demographics include e.g. Fulton et al. (2017); Howard et al. (2012); Mayor et al. (2011); Cassan et al. (2012); Nielsen et al. (2019); Gaia Collaboration et al. (2022); Luque & Pallé (2022). Exoplanet detection campaigns and dedicated missions, such as NASA Kepler, TESS and ESA Cheops, have also significantly contributed to characterise their host stars by measuring key stellar parameters and monitoring their magnetic activity, see e.g. Bastien et al. (2013); Yang et al. (2017); Basri et al. (2013); Balona (2015, 2022); Magrini et al. (2022); Simpson et al. (2023).
Large-scale spectroscopic studies of exoplanetary atmospheres are today in their infancy. Space telescopes, such as NASA Spitzer and NASA/ESA Hubble, have delivered throughout their years of operation large enough data-sets to enable pioneering population studies for tens of exoplanet atmospheres, mainly hydrogen-rich (see e.g., Cowan & Agol (2011); Iyer et al. (2016); Sing et al. (2016); Fu et al. (2017); Tsiaras et al. (2018); Fisher & Heng (2018); Changeat et al. (2022); Edwards et al. (2022)). However, the data quality provided by these space telescopes did not always match the ambition of the scientific community. To enable detailed comparative exoplanetology, population studies should ideally satisfy the following conditions:
Exoplanet spectra should be simultaneously recorded across the broadest possible wavelength range to mitigate the hindrance of observations taken at different epochs (Danielski et al., 2014). This issue is particularly acute if the stellar host is active and/or the atmosphere shows signs of temporal variability due to changes in the cloud coverage, climate or weather patterns (Armstrong et al., 2016; Cohen et al., 2023; Changeat et al., 2024). In particular, optical-to-infrared atmospheric spectra may help to address degeneracies among cloud parameters (Mai & Line, 2019) and to correct for stellar contamination (Ballerini et al., 2012). These mitigations allow to better constrain chemical abundances (Rustamkulov et al., 2023) and the thermal structure of the atmosphere.
@@
While significant strides have been made in characterising exoplanetary atmospheres, less emphasis has been placed so far on using the spectral information in the optical/near-UV regime to detect and correct for potential stellar contamination (e.g. Pinhas et al., 2018; Edwards et al., 2021; Libby-Roberts et al., 2022). The difficulty of this task arises from the diverse ways in which an active star can influence the spectra of its companion planet(s). Stellar magnetic activity results in surface heterogeneities in the form of colder spots and hotter faculae that can cause the transit chord to differ from the observed disk-integrated stellar spectrum. On the stellar photosphere, these heterogeneities can either be occulted or unocculted during the planetary transit w.r.t. our line of sight. While occulted stellar signatures are comparatively easier to identify in transit light curves, unocculted features are trickier to recognise. The first imprint a bump (in the case of a spot) or a depression (for a facula) in the transit light curve, but can be masked or removed by inspection (Narita et al., 2013). On the other hand, unocculted active regions affect the transit depth of the entire light curve, introducing a strong chromatic signal that deepens (decreases) the observed transit depth, overestimating (underestimating) the planetary radius compared to a scenario without spots (faculae) (Rackham et al., 2018). Although stellar contamination effects are strongest in the optical regime, where they introduce characteristic, exponential slopes, they are also non-negligible in the infrared. In this region they are capable of introducing offsets, potentially mimicking or masking absorption features of interest in the planetary atmosphere, e.g. water (Barclay et al., 2021) and/or resulting in erroneous retrieved chemical abundances (Iyer & Line, 2020; Thompson et al., 2024).
@@ 2 Data Analysis
We searched the MAST archive for publicly available exoplanet transit data observed, with at least one grating/grism, by both the STIS and WFC3 instruments on HST. Although we were able to identify more than 30 planets that met this criteria, the data sets considered for this study refer to 20 planets only. To produce a consistent analysis and have a homogeneous data sample, we discarded HAT-P-70 b and KELT-20 b from the final selection as their STIS observations were taken with STIS/G430 at medium resolution. Similarly, we did not include WASP-43 b as its G430L data set did not display a transit and its G750 observation was taken with the medium resolution grating (G750M). WASP-12 b, despite being observed by both grisms on STIS and WFC3, was excluded from our sample due to the contamination by a nearby companion within the system. To obtain a reliable spectrum of WASP-12 b, meticulous decontamination is essential, involving modelling of the companion star. Planetary spectra that displayed strong offsets among their STIS and WFC3 observations were also rejected. These include: HAT-P-17 b, HAT-P-32 b, HAT-P-41 b, HD 189733 b, KELT-7 b, WASP-19 b, WASP-62 b, WASP-76 b, WASP-79 b. Extra processing steps are necessary to correct for the observed offsets and to explain their origin, therefore these planets will be the subject of a separate study. The final sample selection is displayed in Figure 1, while the observations’ proposal ID and the names of their respective principal investigators are reported in Appendix A. There, we also include the instrument-filter combination each planet was observed with and the date each observation was performed. We are aware that some of these exoplanets have also been observed with other instruments, including JWST, but for homogeneity reasons, here we restrict our analysis to data acquired with the two low resolution gratings on STIS and the infrared channel of WFC3. For the avoidance of doubt, each transit spectrum presented here was taken with all or a combination of the G430L and G750L grating on STIS, and the G102 and G141 grism on WFC3, specifically in spatial scanning mode.
While most of the STIS data sets we utilise have undergone prior analysis by different groups employing diverse data reduction pipelines, our study marks the first instance in the literature of analysing the G430L and G750L data sets of HAT-P-18 b and the G430L data set of HAT-P-26 b.
@@
In our wide-sample investigation, we introduce a novel data reduction framework for STIS data, previously presented in Saba et al. (2022), that enables us to uniformly process numerous data sets, with the goal of mitigating biases. WFC3-IR spatial scanned observations were also treated homogeneously across the target sample either by including in our study data sets previously analysed with the Iraclis pipeline or, where new observations were made available, by reducing them with Iraclis (Tsiaras et al., 2016a). Such approach guarantees a standardised methodology for the selected sample and enables a comparative study. Iraclis is a standardised data reduction pipeline that has been extensively used to analyse exoplanet transit and eclipse data. The pipeline was featured in a recent comparison study with two other data analysis pipelines (Mugnai et al., 2024). The authors of the study found that, despite using the same input data and parameters, the output spectra from these pipelines can differ substantially in some cases. An in-depth review of each processing step for each pipeline and their inter-comparison could help identify the origins of these differences and mitigate potential biases. Recently, a multi-pipeline analysis of the WASP-39 b JWST ERS NIRISS data demonstrated that Iraclis produced results consistent with the other pipelines at every stage of the analysis (Feinstein et al., 2023).
@@ 2.2 WFC3 data processing
To ensure homogeneity in the data analysis process, we employed one single pipeline to analyse all transit data sets. Iraclis (Tsiaras et al., 2016a), a specialised WFC3 spatial scanning mode data reduction pipeline and recently adapted to analyse STIS data as well, was used in this study. Most of the WFC3 data sets presented here were previously analysed with Iraclis and collected in two major population studies by Tsiaras et al. (2018) and Edwards et al. (2022). However, these studies focused predominately on WFC3/G141 data, while here we aim to include WFC3/G102 as well. For this reason, we downloaded all publicly available WFC3/G102 data sets for the planets in our sample and processed them with Iraclis. We ensured that the planetary system parameters (Appendix B) we employed were consistent with those used in previous WFC3/G141 studies and across the remaining data sets associated with the same planet. The planets observed with G102 in transmission include HAT-P-11 b, HAT-P-26 b, WASP-17 b and WASP-39 b. These are also the planets that possess a complete spectrum from 0.4 to 1.6 m, i.e. were observed with STIS/G430L, STIS/G750L, WFC3/G102 and WFC3/G141, albeit at different epochs.
@@ 2.3 Light curve fitting
Similarly to WFC3, we attempted to perform the detrending of the STIS spectral light curves with the divide-white method (Kreidberg et al., 2015) but, likely due to chromatic effects arising from the host star and the well-known wavelength-dependent thermal breathing of HST (Brown et al., 2001; Sing et al., 2011), this approach was unsuccessful. A variety of spectroscopic STIS studies have highlighted how the systematics can be both time-dependent (e.g. Nikolov et al., 2014; Lothringer et al., 2018) and wavelength-dependent during a single observation (Huitson et al., 2013), therefore needing a specialised treatment case by case. However, we opted for a uniform systematics model throughout all data sets to maintain consistency: to correct for the combined instrumental and astrophysical trends characteristic of the STIS spectral light curves, we adopted the usual systematics model in Equation 2 and applied it to each spectral light curve separately. Effectively, we modelled the spectral light curves in the same manner as the white light curve. This approach has been used extensively in past STIS studies and has been shown to perform better than employing the divide-white technique (e.g. Chachan et al., 2019). It has consistently proven effective in detrending all data within our sample, thereby ensuring reliable and comparable results throughout.
In the literature there are many examples of how Gaussian Processes (GPs) are employed to fit exoplanet transit light curves (Gordon et al., 2020), especially at the optical wavelengths (Gibson et al., 2017). Generally these algorithms are able to smooth out external features that do not belong to the phenomenon intended to be modelled, in this case by removing any potential spot-crossing signatures from the transit signal. Hence, we attempted to utilise them to detrend STIS light curves as reported in Appendix D. However, we found GPs to be ineffective in providing additional constraints on the light curve modelling. Therefore, all spectral data presented here and employed during the retrievals were produced by applying the standard linear times exponential detrending described in Equation 2.
@@ 2.4 Parameters estimation: and
During the white light curve fitting we fixed the and parameter values to those obtained from literature. Because of the gaps in the light curves, HST observations are not suited to constrain these parameters. A significant coverage in ingress and/or egress is necessary to constrain the semi-major axis and orbital inclination. To test whether combining many HST light curves could be a suitable approach to derive these parameters in the absence of complete light curves, we compared the values obtained by combining three STIS white light curves with those obtained by fitting a complete JWST/NIRISS light curve. We selected WASP-39 b as our candidate planet, as at the time of writing this is the only planet in our sample for which there exist public JWST observations. We used again Equation 2 to detrend the light curves, but leaving and as free parameters. We compared the results obtained using these two approaches with the values listed in ExoClock, and we conclude that the fitting of JWST light curves leads to the most constrained results, which are compatible well within 1 of the ExoClock parameters. The and values obtained from the combined STIS light curve fitting ( = , = ) have a ten times larger uncertainty compared to the JWST fitted values ( = 11.420.03, = 87.770.03, = 0.4440.006) but they are comparable to the ExoClock uncertainties ( = , = ). The semi-major axis and inclination parameters derived from the fitting of the three STIS light curves are 2 and 3 away respectively from both JWST and ExoClock, while JWST and ExoClock parameters agree within 1. Therefore, rather than trying to constrain these orbital parameters by fitting multiple incomplete light curves together, we suggest employing either tabulated parameters or fitted parameters from complete white light curves and fixing them during the spectral light curve fitting. For consistency we fitted all white and chromatic light curves using the ExoClock parameters as fixed values across the whole planetary sample, including on WASP-39 b.
@@ 3 Spectral Modelling
Due to the often divergent STIS spectra resulting from their acquisition at varying epochs (Fig. 2), we decided to model the different combinations of data sets as separate spectral scenarios for each planet. As reported in Table 1, each of the potential combinations of data sets is identified by a Case number. The spectral modelling is performed with the fully-Bayesian retrieval framework TauREx 3.1 (Al-Refaie et al., 2021, 2022). The software has been utilised in a variety of modelling comparison projects (Barstow et al., 2020, 2022; Powell et al., 2024; Villanueva et al., 2024), producing results that align well (within 1) with other retrieval tools. The robustness and consistency of TauREx 3.1 highlights its reliability as an atmospheric retrieval tool. Through the use of nested sampling techniques (Skilling, 2006), the retrieval code maps the correlations among the atmospheric parameters in the defined parameter space. TauREx operates on an entirely open-source framework. Utilising the plug-in system, users have the flexibility to integrate their own models, allowing for the representation of various atmospheric conditions such as stellar contamination, cloud parameterisation, and temperature-pressure profiles of their choosing. To explore the parameter space, we employed the widely utilised nested sampling algorithm MultiNest (Feroz et al., 2009; Buchner et al., 2014) with 250 live points and an evidence tolerance set to 0.5.